Shutdown pquest SNF 2011-02-18 10:26:30: chamber clean and pm

eenriquez at eenriquez at
Fri Feb 18 15:01:16 PST 2011

Completed chamber wet clean and monthly pm.  Replaced damaged cooling helium o-ring.  Baked out the new o-ring for 3 hours at 190 C.  Cycled wafers with no handling problem.  The cooling He flow is high at 7.4.  Need to level the clamp.

More information about the pquest-pcs mailing list